OpenStudy (anonymous):

x^2-4/x^2-2x I need help finding the vertical and horizontal asymtptotes and what kind of symmetry if yall dont mind :D

OpenStudy (anonymous):

Is the function : \[f(x)=x^2-\frac4{x^2}-2x???\]

OpenStudy (anonymous):

no its (x^2-4) divided by (x^2-2x)

OpenStudy (anonymous):

LIke this : \[f(x)=\frac{x^2-4}{x^2-2x}??\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

First, you have to determine the domain of the function, what is it ?

OpenStudy (anonymous):

the domain is (negative infinity, 0)(0,2)(2,infinity)

OpenStudy (anonymous):

Exactly : The domaine is : \[D=(-\infty, 0)\cup(0,2)\cup (2,+\infty)\] So you have to calculate the limits at the bounds of D

OpenStudy (anonymous):

yo lost me haha how do i do that

OpenStudy (anonymous):

It is very simple, let's begin with the infinity : \[\Large \lim_{x\to+\infty}f(x)=\lim_{x\to+\infty}\frac{x^2-4}{x^2-2x}=\lim_{x\to+\infty}\frac{x^2}{x^2}=\lim_{x\to+\infty}1=1\] and : \[\Large \lim_{x\to-\infty}f(x)=\lim_{x\to-\infty}\frac{x^2-4}{x^2-2x}=\lim_{x\to-\infty}\frac{x^2}{x^2}=\lim_{x\to-\infty}1=1\] So we can say that the horizontal line with equation (y=1) is an horizontal asymptote in the neighborhood of +infty and -infty

OpenStudy (anonymous):

i see okay what about the vertical asymptote

OpenStudy (anonymous):

Now we calculate the limit at 0 from the two sides : \[\Large \lim_{x\to0^-}f(x)=\frac{0^2-4}{0^+}=-\infty\] and : \[\Large \lim_{x\to0^+}f(x)=\frac{0^2-4}{0^-}=+\infty\] So the line with equation (x=0) is a vertical asymptote. the same thing with 2 :P

OpenStudy (anonymous):

so the vertical asymptote is 0

OpenStudy (anonymous):

A vertical asymptot is a line not a number O.o After that: this function has 2 verticals asymptotes ;)

OpenStudy (anonymous):

where would i plot the vertical asymptotes i mean. sorry!

OpenStudy (anonymous):

at 0 and where else

OpenStudy (anonymous):

and 2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay one last thing! what about the symmetry how do i figure out if its symmetrical. it has something to do with f(-x)=f(x) but im not clear on it